# Cost Functions in Economic Complexity

### v2

@@ Acknowledgments:
Conflicts of interest:
The authors declare no conflict of interest.
Abbreviations:
The following abbreviations are used in this manuscript:
@@ A.1 ECI
$\dfrac{\partial\dot{Z_{i}}}{\partial Z_{h}}=\dfrac{\partial\dot{Z_{h}}}{\partial Z_{i}}$ (15)
Given that the full adjacency matrix $A_{ij}$ is symmetric by construction, both quantities are equal to $\dfrac{A_{ih}}{\sqrt{k_{i}k_{h}}}+\delta_{ih}$.
@@ Appendix B EFC: Solution in the case N=2
$\begin{cases}V^{(n+1)}_{1}=\delta+\frac{1}{V^{(n)}_{2}},\\ V^{(n+1)}_{2}=\delta+\frac{1}{V^{(n)}_{1}}\end{cases}$
@@ Appendix C EFC: Regular graphs
$V^{*}=\sqrt{K+\tfrac{1}{4}\delta^{2}}+\delta.$
To further formalize this convergence, denote by $X^{(n)}_{m}$ and $Y^{(n)}_{m}$ the orbits starting from initial values $X^{(0)}_{m}\leq V^{(0)}_{m}$ and $Y^{(0)}_{m}\leq V^{(1)}_{m}$, respectively. Then, by the inequalities in (22) and the monotonicity of $f_{K}$,
@@
In summary, the sequences $\{V^{(n)}_{m}\}$ and $\{V^{(n)}_{M}\}$ are bounded respectively from below and above by trajectories of the dynamical system defined in (23). Consequently, we obtain
@@
It is worthwhile to note that, from (22) and the lower bound in (21), one deduces that
$V^{(n)}_{M}-V^{(n)}_{m}\leq\frac{K^{2}\left(V^{(n-2)}_{M}-V^{(n-2)}_{m}\right)}{(\delta V^{(n-2)}_{M}+K)(\delta V^{(n-2)}_{m}+K)}\leq\frac{K^{2}}{(K+\delta^{2})^{2}}\left(V^{(n-2)}_{M}-V^{(n-2)}_{m}\right)$.
Since $\frac{K^{2}}{(K+\delta^{2})^{2}}<1$, this inequality shows that the convergence in (24) is exponentially fast in time. Moreover, applying (24) to the stationary case, we deduce that the fixed point equation
@@
As discussed at the end of Subsection 3.2.2 in the explicit case $N=2$, the presence of the positive parameter $\delta$ is crucial for ensuring convergence to a unique fixed point. When $N>2$ and $\delta=0$, the argument leading to (24) fails because the function $f_{K}(x)$ becomes the identity $f_{K}(x)=x$. Moreover, unlike the special case $N=2$, it is not even clear whether $\vec{V}^{*}=(\sqrt{K},\ldots,\sqrt{K})$ is the unique fixed point.
@@
$\frac{\partial U}{\partial V_{i}}=-\frac{1}{V_{i}^{2}}\sum_{j}\frac{A_{ij}}{V_{j}}+\frac{1}{V_{i}}-\frac{\delta}{V_{i}^{2}}=-\frac{1}{V_{i}^{2}}\left(\sum_{j}\frac{A_{ij}}{V_{j}}-V_{i}+\delta\right).$
Hence, the condition for a stationary point, $\nabla U(\vec{V})=0$, is equivalent to the fixed point equation (25). Since $U(\vec{V})$ diverges to $\infty$ as $\vec{V}$ approaches the boundary of $(0,+\infty)^{N}$, we conclude that the fixed point $\vec{V}^{*}$ is indeed the unique absolute minimizer of $U$ in $(0,+\infty)^{N}$.
@@ Appendix D EFC: local convergence for general graphs
Drawing upon the results established in Section 3.2.1, the mapping $\vec{F}$ possesses precisely one fixed point, which coincides with the unique minimizer $\vec{V}^{*}$ of the cost function $U$. We aim to demonstrate that $W=U-\min U$ serves as a Lyapunov function for the equilibrium $\vec{V}^{*}$. According to standard stability theory, this would imply that the equilibrium is asymptotically stable.
@@
$\begin{split}U(F(\vec{V}))-U(\vec{V})&=\nabla U(\vec{V})\cdot\vec{D}+\int_{0}^{1}\!\mathrm{d}\lambda\,\lambda\vec{D}\cdot D^{2}U(\vec{V}+\lambda\vec{D})\vec{D}\\ &=-\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}+\frac{1}{2}\vec{D}\cdot D^{2}U(\vec{V}^{*})\vec{D}+\mathcal{O}(\varepsilon)|\vec{D}|^{2}\qquad\forall\,\vec{V}\in B,\end{split}$
@@
$\begin{split}\frac{1}{2}\sum_{i,j}\frac{A_{ij}D_{i}D_{j}}{(V_{i}^{*})^{2}(V_{j}^{*})^{2}}&=\frac{1}{2}\sum_{i,j}\frac{A_{ij}}{V_{i}^{*}V_{j}^{*}}\frac{D_{i}}{V_{i}^{*}}\frac{D_{j}}{V_{j}^{*}}\leq\frac{1}{4}\sum_{i,j}\frac{A_{ij}}{V_{i}^{*}V_{j}^{*}}\left(\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}+\frac{D_{j}^{2}}{(V_{j}^{*})^{2}}\right)\\ &=\frac{1}{2}\sum_{i,j}\frac{A_{ij}}{V_{i}^{*}V_{j}^{*}}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}\\ &=\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}\sum_{j}\frac{A_{ij}}{V_{j}^{*}}=\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}(V_{i}^{*}-\delta)\\ &=\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}-\frac{\delta}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}\end{split}$
@@
$U(F(\vec{V}))-U(\vec{V})\leq-\frac{\delta}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}+\mathcal{O}(\varepsilon)|\vec{D}|^{2}\leq\left(-\frac{\delta}{2(V_{M}^{*})^{3}}+\mathcal{O}(\varepsilon)\right)|\vec{D}|^{2}\qquad\forall\,\vec{V}\in B$

### v3

@@ A.1 ECI
[…]
@@ Appendix B EFC: Solution in the case N=2
$\cases{V}^{(n+1)}_{1}=\delta+\frac{1}{V^{(n)}_{2}},\\ V^{(n+1)}_{2}=\delta+\frac{1}{V^{(n)}_{1}}$
@@ Appendix C EFC: Regular graphs
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\frac{\partial U}{\partial V_{i}}=-\frac{1}{V_{i}^{2}}\sum_{j}\frac{A_{ij}}{V_{j}}+\frac{1}{V_{i}}-\frac{\delta}{V_{i}^{2}}=$
$\frac{1}{V_{i}^{2}}\left(\sum_{j}\frac{A_{ij}}{V_{j}}-V_{i}+\delta\right).$
[…]
@@ Appendix D EFC: local convergence for general graphs
Drawing upon the results established in Section III.2.1, the mapping $\vec{F}$ possesses precisely one fixed point, which coincides with the unique minimizer $\vec{V}^{*}$ of the cost function $U$. We aim to demonstrate that $W=U-\min U$ serves as a Lyapunov function for the equilibrium $\vec{V}^{*}$. According to standard stability theory, this would imply that the equilibrium is asymptotically stable.
@@
[…]
@@
[…]
@@
[…]